# villa-X: Enhancing Latent Action Modeling in Vision-Language-Action Models

### v2

@@ A.2 Actor Module
We extend our policy head with a variant of HPT [62], assigning each embodiment its own pair of state‐ and action‐projection layers while sharing all other parameters. Visual features from the wrist camera are extracted by a pretrained ResNet-18 [24] and fused into the main model via a shared cross-attention head that maps the ResNet features into $16$ tokens. During training, wrist-view inputs are randomly masked 50% of the time. We also observed that the latent‐action representation can be overly exploited by the robot‐action expert, so we regularize this with two complementary dropout schemes. First, we add a 50% attention-weight dropout on the latent‐action stream. For the remaining tokens, we randomly mask 50% latent action tokens. This combined masking strategy encourages the model to learn robust, generalizable policy that will balance the predicted latent actions as well as the input image and instruction. Each expert contains approximately 300 M parameters and is trained from scratch. We train all components jointly using a learning rate of 5e55e-5 with a $200$-step linear warmup. We clip gradients to a maximum norm of 1.0 to ensure stable optimization. The pretraining takes 4 days on 64 NVIDIA A100 GPUs.
Appendix B Dataset
B.1 Data Mixture
We curated a data mixture by combining both robot data and action-free human videos for our pretraining phase. For robot data, we draw primarily from OpenX [13] mixture and AgiBot [1]. For OpenX dataset, our base data mixture is created primarily based on [30, 54]. In total, we use 223.5M trajectories with 1.6M frames of robot data. For human videos, we use a mixture of Ego4D [22], EgoPAT3D [37], EGTEA Gaze+ [36], EPIC-KITCHENS [16], HO-Cap [61], HOI4D [44], HoloAssist [64], RH20T [19], Something Something V2 [21]. Altogether, this yields 3.6M clips of human videos. During LAM pretraining, we exclusively utilize the primary third-person camera view. For policy pretraining, we optionally incorporate the wrist-mounted view (when available), applying a 5050% dropout. A full breakdown of our data mixture is listed in Table 6.
B.2 Data Preprocessing
@@
Appendix C LAM visualization results
C.1 Image Pairs with Similar Latent Actions
@@
Figure 9 visualizes additional image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors.
@@
Appendix D Simulation Evaluation Details
D.1 SIMPLER Benchmark
@@
D.2 LIBERO
The LIBERO benchmark [40] evaluates knowledge transfer in multitask and lifelong robot learning problems for robotic manipulation, consisting of four task suites: LIBERO-Spatial, LIBERO-Goal, LIBERO-Object, LIBERO-Long. For fine-tuning of our models, we reinitialize the linear state encoder, action encoder, and action decoder, and tune the full parameters (except for the vision encoder). We fine-tune all the models for 50k gradient steps on the fine-tuning dataset on each task suite of LIBERO. We follow the standard evaluation protocol to test by randomizing the initial states of the environments. Reported results (Table 3) are the average success rates across on a total of 200 trials and a total of 20 trials per task on each task suite. Please refer to Section 4.2 and Appendix C of LIBERO [40] for more details.
Appendix E Real-world Robot Platforms Evaluation Details
E.1 Realman robot arm
@@
E.2 XHand dexterous hand

### v3

@@ Appendix B Additional Implementation Details for Actor Module
We extend our policy head with a variant of HPT [63], assigning each embodiment its own pair of state‐ and action‐projection layers while sharing all other parameters. Visual features from the wrist camera are extracted by a pretrained ResNet-18 [23] and fused into the main model via a shared cross-attention head that maps the ResNet features into $16$ tokens. During training, wrist-view inputs are randomly masked 50% of the time. We also observed that the latent‐action representation can be overly exploited by the robot‐action expert, so we regularize this with two complementary dropout schemes. First, we add a 50% attention-weight dropout on the latent‐action stream. For the remaining tokens, we randomly mask 50% latent action tokens. This combined masking strategy encourages the model to learn robust, generalizable policy that will balance the predicted latent actions as well as the input image and instruction. During training, we sample $\tau$ from different beta distributions for latent actions and robot actions, which biases the timesteps for latent actions towards the noisier regime. Each expert contains approximately 300 M parameters and is trained from scratch. We train all components jointly using a learning rate of 5e55e-5 with a $200$-step linear warmup. We clip gradients to a maximum norm of 1.0 to ensure stable optimization. The pretraining takes 4 days on 64 NVIDIA A100 GPUs.
Appendix C Dataset
C.1 Data Mixture
We curated a data mixture by combining both robot data and action-free human videos for our pretraining phase. For robot data, we draw primarily from OpenX [12] mixture and AgiBot [1]. For OpenX dataset, our base data mixture is created primarily based on [29, 52]. In total, we use 1.6M trajectories with 223.5M frames of robot data. For human videos, we use a mixture of Ego4D [21], EgoPAT3D [37], EGTEA Gaze+ [36], EPIC-KITCHENS [15], HO-Cap [62], HOI4D [42], HoloAssist [64], RH20T [18], Something Something V2 [20]. Altogether, this yields 3.6M clips of human videos. During LAM pretraining, we exclusively utilize the primary third-person camera view. For policy pretraining, we optionally incorporate the wrist-mounted view (when available), applying a 5050% dropout. A full breakdown of our data mixture is listed in Table 5.
C.2 Data Preprocessing
@@
Appendix D LAM visualization and More Ablations
D.1 Image Pairs with Similar Latent Actions
@@
Figure 6 visualizes image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors.
@@
D.2 Transfer Video Demonstrations into Robot Actions through LAM and Proprio FDM
Figure 7: Transfer robot video demonstrations into robot actions through LAM and proprio FDM in SIMPLER simulator. Upper: the SIMPLER rollout closely reproduce the motion of moving downwards, Bottom: the SIMPLER rollout closely reproduce the motion of moving right.
Figure 8: Transfer human video demonstrations into robot actions through LAM and proprio FDM in SIMPLER simulator. Upper: the SIMPLER rollout closely reproduce the motion of moving right; Middle: the SIMPLER rollout closely reproduce the motion of moving forward and backward; Bottom: the SIMPLER rollout closely reproduce the motion of moving right.
To further demonstrate the transfer ability of our LAM, we extract latent actions from videos of task demonstrations, map them to robot actions using the proprio FDM, and execute the resulting robot actions in the SIMPLER simulator.
The results are presented in Figure 7 and Figure 8. In each figure, the top row shows the video demonstrations used by LAM to extract latent actions, while the bottom row displays the corresponding SIMPLER simulation results, where real actions decoded from the latent actions using proprioceptive FDM are executed. Specifically, Figure 7 illustrates robot-to-robot transfer, and Figure 8 illustrates human-to-robot transfer. The simulated motions closely reproduce the original demonstrations, indicating that latent actions learned by villa-X are both aligned with and grounded in the robot’s actions.
D.3 More Ablations on LAM
To validate the contribution of the embodiment context in our proprio-FDM, we further conducted an ablation study comparing our full method (”Ours”) against a version without the context (”Ours w/o context”). Both models were trained on 10
(1) Performance on validation dataset: We measured the reconstruction loss of visual FDM and proprio FDM on the validation set:
Method Visual FDM loss ($\downarrow$) Proprio FDM loss ($\downarrow$) Ours w/o context 0.068 0.078 Ours 0.057 0.070 Relative improvement 16.2% 10.3%
Table 6: Performance comparison on the validation set.
(2) Zero-Shot Generalization to a Novel Embodiment: We evaluated the model on our dataset collected on our Realman robot arm dataset (from Section 4.4), an embodiment completely unseen during training. We then conducted the action probing experiment described in Section 4.1 by inferring latent actions with IDM and training a new MLP to predict robot actions from the latent actions.
Method Probing loss ($\downarrow$) Probing loss (xyz) ($\downarrow$) Probing loss (rot) ($\downarrow$) Probing loss (gripper) ($\downarrow$) Ours w/o context 0.165 0.0675 0.00861 0.928 Ours 0.152 0.0574 0.00619 0.873 Relative improvement 7.9% 15.0% 28.1% 5.9%
Table 7: Zero-shot generalization to an unseen embodiment.
The results from both experiments demonstrate that the embodiment context improves performance and aids generalization to novel embodiments. We hypothesize that while the visual FDM provides general transferability by aligning latent actions with visual changes, the proprio-FDM grounds these latent actions in robot physical dynamics. However, due to data heterogeneity (e.g., different action definitions / controllers, as discussed previously), the model requires the embodiment context to disambiguate different embodiments and learn a more consistent, grounded latent action space.
Appendix E Latent Action Expert Visualization
Figure 9: Generated image sequence jointly by the latent expert and the world model via latent actions, following different instructions from the same initial image (Part I).
Figure 10: Generated image sequence jointly by the latent expert and the world model via latent actions, following different instructions from the same initial image (Part II).
In this experiment, we demonstrate the performance of the latent expert by passing its prediction through the image reconstruction FDM that takes the latent action as inputs and predicts the future observations, which forms a simulated environment for the iteratively executing the latent expert.
Starting from a single initial image, the latent expert and image reconstruction FDM jointly generate different behaviors in videos that follow diverse instructions using only latent actions. We experiment with initial images from RT-1 and Bridge dataset, and show the image clips of generated videos in Figure 9 and Figure 10 with different language instructions. The results show that the latent expert properly follows the language instructions for task solving, where the latent expert properly recognizes the target objects and predict latent actions that move towards the target object.
Appendix F More Ablations on policy model
We primarily conducted ablation studies on two main components: (1) the attention mask and (2) the embodiment context. Our experiments follow the same setting as Table 1 in the main paper. The ablation results below show that both the attention mask and embodiment context are effective in improving performance on two robot platforms: Google Robot and WidowX Robot.
Table 8: Ablation study results for the policy model. The first columns (Pick, Move, Drawer, Place) refer to the Google Robot, and the last columns (Carrot, Eggplant, Spoon, Cube) refer to the WidowX Robot. All numbers are success rates (%).
Method Pick Move Drawer Avg. Carrot Eggplant Spoon Cube Avg. Ours 81.7 55.4 38.4 58.5 24.2 71.7 48.3 19.2 40.8 Ours w/o mask 80.3 30.6 48.8 53.2 18.3 52.5 38.3 26.7 34.0 Ours w/o context 86.6 21.3 39.3 49.1 28.3 67.5 25.8 32.5 38.5
Appendix G Simulation Evaluation Details
G.1 SIMPLER Benchmark
@@
Appendix H LIBERO Benchmark
Baselines and Experimental Setup
We compare with the following existing models: Diffusion Policy [11] trained from scratch, Octo [52], OpenVLA [29], $\pi_{0}$ [4], $\pi_{0}$ FAST [54], TraceVLA [71] and SpatialVLA  [55]. For $\pi_{0}$, we use the open source version [57] and the same training set as our model. All models follow a two-stage pretraining-finetuning protocol. We finetune villa-X and villa-X w/o latent on the demonstration data of the each task suite separately, and test on the LIBERO simulator for 10 tasks and 20 trials per task on each task suite.
Experimental Results
Table 9 summarizes the success rates on each task suite of LIBERO. Our model achieves better performance than existing methods in all the four task suites. Also, our model with latent action achieves higher performance on all the four task suites and average performance, confirming that the proposed latent action expert improves the manipulation performance.
Method Spatial Object Goal Long Average Diffusion Policy [11] 78.3 92.5 68.3 50.5 72.4 Octo-base [52] 78.9 85.7 84.6 51.1 75.1 OpenVLA [29] 84.7 88.4 79.2 53.7 76.5 $\pi_{0}$ (reimplement [57]) 88.0 88.5 87.0 61.0 81.1 $\pi_{0}$-FAST [54] 96.4 96.8 88.6 60.2 85.5 TraceVLA [71] 84.6 85.2 75.1 54.1 74.8 SpatialVLA  [55] 88.2 89.9 78.6 55.5 78.1 Ours w/o latent 86.0 86.5 85.0 70.0 81.9 Ours 97.5 97.0 91.5 74.5 90.1
Appendix I Real-world Robot Platforms Evaluation Details
I.1 Realman robot arm
@@
I.2 XHand dexterous hand